i) If Rs.64 amount to Rs.83.20 in 2 years, what will Rs.86 amount to in 4 years at the same rate 

of simple interest?


1
Expert's answer
2021-05-27T12:19:17-0400

Simple Interest


"I=Prt"

"I=" Interest

"P=" Principal (Initial Value)

"r=" Interest Rate

"t=" time (years)

"r=" const

If Rs.64 amount to Rs.83.20 in 2 years


"P=Rs.64, t=2\ years"


"Rs.83.20-Rs.64=Rs.19.20"

"r=\dfrac{Rs.19.20}{Rs.64\times2}=0.15 (15\%)"



Given "P_{new}=Rs.86, r=0.15, t_{new}=4\ years"


"I_{new}=P_{new}rt_{new}"

"I_{new}=Rs.86(0.15)(4)=Rs.51.60"

"A_{new}=P_{new}+I_{new}"

"=Rs.86.00+Rs.51.60=Rs.137.60"

"Rs.86" amount amount to in 4 years at the same rate of simple interest will be "Rs.137.60."